# BokehDiff: Neural Lens Blur with One-Step Diffusion

### v1

@@ Abstract
$\dagger$\$$\dagger$\$footnotetext: Work done during an internship at Vivo.$\ast$\$$\ast$\$footnotetext: Corresponding authors.
{strip}[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Figure 1: BokehDiff bridges the gap between physics and diffusion priors, and is able to synthesize photorealistic lens blur effects even when inaccurate depth estimation causes previous methods (BokehMe [38], MPIB [39], and Dr. Bokeh [47]) to fail, especially at the depth discontinuities. The examples show previous methods over-blur the horse’s tail, the person’s hair, and the whiskers of the cat.
@@ 1 Introduction
The bokeh effect is the out-of-focus blurriness observed in photos, physically caused by using a lens with a large aperture, and is often used in portrait photography to emphasize the subject. Due to the cost of large aperture lenses, bokeh rendering has become a hot topic in the computational photography community. Previous works [39, 38, 28, 66] mostly aim to simulate the blurriness accurately with a pixel-level accurate depth estimation. However, since depth prediction tends to fail on edges and intricate details, artifacts can often be observed on structures such as people’s hair and animals’ fur, as shown in Fig. 1. As state-of-the-art diffusion models (e.g., SDXL [42]) are already capable of generating photorealistic lens blur effects from text instructions [61], can they be applied to render lens blur effects from a given image?
@@ 2.1 Bokeh Rendering
Neural rendering uses a neural network to mimic the image formation model, and is often trained end-to-end on synthetic data with ground truth depth map [38, 53, 28, 20, 44, 13, 60, 33]. As the network learns to add specific amount of blur to the input image, the problem of inaccurate depth estimation constitutes the major bottleneck of performance, as seen in Fig. 1.
@@ 3.2 Physics-Inspired Self-Attention Module
$C_{qk}=\text{Soft}[r_{c}(k)-c_{i}\cdot\|P_{q}-P_{q}\|_{2}].$ (6)
@@ 3.4 Supervision
$\mathcal{L}_{\text{D}}=\mathbb{E}_{I}[\log D(I_{b})]+\mathbb{E}_{\hat{I}}[\log(1-D(\hat{I}))],$ (11)
while $\mathcal{L}_{\text{adv}}=-\mathbb{E}_{\hat{I}}[\log D(\hat{I})]$ is used for finetuning diffusion model. In all, the finetuning loss is given by

### v2

@@ Abstract
[…]
@@ 1 Introduction
The bokeh effect is the out-of-focus blurriness observed in photos, physically caused by using a lens with a large aperture, and is often used in portrait photography to emphasize the subject. Due to the cost of large aperture lenses, bokeh rendering has become a hot topic in the computational photography community. Previous works [39, 38, 28, 66] mostly aim to simulate the blurriness accurately with a pixel-level accurate depth estimation. However, since depth prediction tends to fail on edges and intricate details, artifacts can often be observed on structures such as people’s hair and animals’ fur, as shown in LABEL:fig:teaser. As state-of-the-art diffusion models (e.g., SDXL [42]) are already capable of generating photorealistic lens blur effects from text instructions [61], can they be applied to render lens blur effects from a given image?
@@ 2.1 Bokeh Rendering
Neural rendering uses a neural network to mimic the image formation model, and is often trained end-to-end on synthetic data with ground truth depth map [38, 53, 28, 20, 44, 13, 60, 33]. As the network learns to add specific amount of blur to the input image, the problem of inaccurate depth estimation constitutes the major bottleneck of performance, as seen in LABEL:fig:teaser.
@@ 3.2 Physics-Inspired Self-Attention Module
$C_{qk}=\text{Soft}[r_{c}(k)-c_{i}\cdot\|P_{q}-P_{k}\|_{2}].$ (6)
@@ 3.4 Supervision
$)],\mathcal{L}_{\text{D}}=\mathbb{E}_{I}[\log D(I_{b})]+\mathbb{E}_{\hat{I}}[\log(1-D(\hat{I}$))], (11)
[…]